Compute the first and second order derivatives of the following:
1. Y = lnx + 3x – 2
2. Y = x
1
Expert's answer
2021-08-03T12:31:26-0400

"y=lnx+3x-2"

First derivative

"y'=\frac {d}{dx}[lnx+3x-2]"


"y'=\frac {d}{dx}[lnx]+3\frac{d}{dx}[x]+\frac{d}{dx}[-2]]"


"y'=\frac {1}{x}+3(1)+0"


"y'=\frac {1}{x}+3" "....................1st"


Second derivative

"y''=\frac{d}{dx}[\frac {1}{x}+3]"


"y'=-\frac {1}{x^2}" "....................2nd"




Then for


"y=x"

"y'=\frac{d}{dx}[x]=1" ........1st derivative


"y''=\frac{d}{dx}[1]= 0" .........2nd derivative